# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v1

@@ 6.1 Noncooperation with ETS under the Baseline Emission Caps
Figure 2 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2032, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 2 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
Global emissions are constrained by the global emission caps starting from 2032. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volumes of both emission abatement and traded permits, along with their associated costs. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.2 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$749 per ton of carbon in 2050 and \$2,225 in 2069 (Figure 2 bottom-left). The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions. As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050 for the US, EU, Japan, and OHI regions, and 2060 for Russia and China. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^8] In the bottom-left panel of Figure 2, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.74 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right). Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 3 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage. After 2032, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Middle East, Africa, and OHI regions, with the US and India being the largest permit buyers. The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider. Japan is expected to be involved in a relatively small volume of permit trading.
@@
In Figure 4, the MAC of each region gradually decreases below the permit price after its net zero emissions (before trading of permits) is achieved. Russia is the first region to achieve net zero emissions (before trading of permits) in 2051 (as shown in Figure A.6 in Appendix A.8.1). But Russia still sells permits afterwards, as shown in Figure 3, because its after-trade net zero emission target year is 2060 under the baseline emission cap scenario. After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound. Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China and Eurasia) and the 2060s (the US, the EU, Japan, MidEast, Latin America, and the OHI). India, Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. [^10] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Recall that, when the regional net zero emissions (before trading of permits) are achieved, the optimal carbon tax should be larger than the gap between MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding, then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
@@
We further confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4. For example, in 2050, the MAC for the US is \$908 per ton of carbon, the permit price is \$749, and their difference is exactly equal to the regional SCC of the US, \$159. Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC, starting at \$105 per ton of carbon in 2020 and steadily increasing up to \$210 per ton of carbon by the end of this century. The US is followed by the EU (from \$42 to \$138 per ton of carbon) and China (from \$39 to \$115 per ton of carbon) in this century. The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon). Our results show that there is considerable heterogeneity in the SCC across regions and suggest that a significant amount of regional carbon tax implementation is required, particularly for the US, the EU, and China, under the global ETS with the baseline emission cap scenario.
@@ 6.2 Comparative Analysis of Different Cases
We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation. Specifically, cooperation results in approximately 0.19 degrees Celsius lower temperature with the ETS and 0.15 degrees Celsius lower temperature without the ETS. Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps.
@@
When comparing the cases with and without the ETS, we find that the ETS may play different roles under noncooperation and cooperation. In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS. This is because with the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level. This happens until 2050. In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time. With the ETS, slightly higher levels of global emissions and temperature rise are observed before 2040 and 2050, respectively, and the opposite trends are observed after that. This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
@@
We also compare the regional MAC and SCC across the different cases. Figure 7 presents the regional MAC and SCC under both the cooperative and the noncooperative models with or without the ETS, using the US as an illustrative example. [^11] For comparisons of other regions, refer to Appendices A.8.4 and A.8.3. Under cooperation, the MAC of the US (Figure 7 top-left) is very large in the initial years. This is because the US, as a developed region, needs to mitigate a significant amount of emissions so that developing regions can mitigate less (as shown in Figure A.9 in Appendix A.8.3), thus improving global social welfare. After the 2030s, however, the US net emissions become zero in the cooperative case [^12] For regional net emissions under cooperation and noncooperation, see Appendix A.8.2. and cannot be further mitigated, so its MAC starts to decline over time. The effect of the ETS depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU at around 2040); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm).
@@
The top-right panel of Figure 7 shows that the cooperative SCC of the US is sustained at a high level throughout all periods. This is consistent with the high MAC of the US. When the US regional net emissions are not binding at zero before 2032 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2032. Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions. This also implies that a developed region will contribute more to mitigating emissions under cooperation. This pattern is consistent across regions: Figure A.11 in Appendix A.8.4 shows that the US, EU, Japan and OHI have a high cooperative SCC, while the developing regions like India and Africa have a small cooperative SCC.
Under noncooperation, the MAC of the US is substantially lower in the case with the ETS compared to the case without the ETS (Figure 7, bottom-left panel). Such discrepancy persists until 2058, the period during which the US is a net permit buyer (shown in Figure 3), reflecting the fact that implementation of the ETS offers the US a more cost-effective means of keeping its net emissions below its cap. Consequently, the US conducts less emission abatement. For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2032 (see Figure A.10 in Appendix A.8.3).
@@
Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS. For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2032 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2032 (see the two top panels of Figure 7), which is equal to the optimal regional carbon tax under cooperation. For the noncooperative case without the ETS, the US regional net emissions are binding at their caps (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC (see the two bottom panels of Figure 7), which is equal to the optimal regional carbon tax under noncooperation. Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Russian under cooperation until 2057 in Figures A.7, A.9, A.11); and Russian under noncooperation until 2042 in Figures A.8, A.10, and A.12). This relationship holds for every region as shown in Figures A.7-A.12.
The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under both cooperation and noncooperation. This pattern holds for every region as shown in Figures A.11 and A.12.
@@ 6.3 Alternative Net Zero Scenarios
Along with the baseline emission cap scenario, we further analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090. The top-left and top-right panels in Figure 8 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,538 per ton of carbon in 2049, and the temperature rise is restricted to 1.64 degrees Celsius by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$344 and \$158 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.86 degrees Celsius and 2.08 degrees Celsius above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
@@
The bottom-left and bottom-right panels in Figure 8 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.8.5, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2086 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$216 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$193 in 2100 in the net zero 2090 scenario. This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.

### v2

@@ 6.1 Noncooperation with ETS under the Baseline Emission Caps
Figure 2 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2030, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 2 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
Global emissions are constrained by the global emission caps starting from 2030. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volumes of both emission abatement and traded permits, along with their associated costs. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.18 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$811 per ton of carbon in 2050 and \$2,254 in 2069 (Figure 2 bottom-left). The kink in the permit price path in 2050 is a result of some regions achieving net zero emissions. As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050s for Japan (2050) and EU (2053), 2060 for Russia and China, and 2070 for the rest of regions. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^10] In the bottom-left panel of Figure 2, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right), restricted by the emission caps. [^11] If there is no emission cap, then the temperature anomaly will be much higher. This can be seen in a later discussion with alternative emission caps. Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 3 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage. After 2029, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Africa, and OHI regions, with the US, EU, and India being the largest permit buyers. The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider. Japan is expected to be involved in a relatively small volume of permit trading.
@@
In Figure 4, the MAC of each region gradually decreases below the permit price after its net zero emissions (before trading of permits) is achieved. Russia is the first region to achieve net zero emissions (before trading of permits) in 2051 (as shown in Figure A.6 in Appendix A.8.1). But Russia still sells permits afterwards, as shown in Figure 3, because its after-trade net zero emission target year is 2060 under the baseline emission cap scenario. After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound. Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China, Latin America, and the OHI) and the 2060s (the US, the EU, Japan, Eurasia, India, and MidEast). Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. [^13] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Recall that, when the regional net zero emissions (before trading of permits) are achieved, the optimal carbon tax should be larger than the gap between MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding, then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
@@
We further confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4. For example, in 2050, the MAC for the US is \$971 per ton of carbon, the permit price is \$811, and their difference is exactly equal to the regional SCC of the US, \$160. Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC, starting at \$106 per ton of carbon in 2020 and steadily increasing up to \$213 per ton of carbon by the end of this century. The US is followed by the EU (from \$42 to \$139 per ton of carbon) and China (from \$39 to \$117 per ton of carbon) in this century. The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon). Our results show that there is considerable heterogeneity in the SCC across regions and suggest that a significant amount of regional carbon tax implementation is required, particularly for the US, the EU, and China, under the global ETS with the baseline emission cap scenario.
@@ 6.2 Comparative Analysis of Different Cases
We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation. Specifically, cooperation results in approximately 0.17 degrees Celsius lower temperature with the ETS and 0.13 degrees Celsius lower temperature without the ETS. Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps. Moreover, the ETS has little impact on the temperature path, by reducing the temperature anomaly in 2010 with only 0.02 degrees Celsius, as the emission caps exist no matter whether the ETS exists or not.
@@
When comparing the cases with and without the ETS, [^14] Note that the emission caps are still there, no matter whether the ETS is implemented or not. we find that the ETS may play different roles under noncooperation and cooperation. In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS. This is because with the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level. This happens until 2050. In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time. With the ETS, lower levels of global emissions are observed after 2030. This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
@@
We also compare the regional MAC and SCC across the different cases. Figure 7 presents the regional MAC and SCC under both the cooperative and the noncooperative models with or without the ETS, using the US as an illustrative example. [^15] For comparisons of other regions, refer to Appendices A.8.4 and A.8.3. Under cooperation, the MAC of the US (Figure 7 top-left) is very large in the initial years. This is because the US, as a developed region, needs to mitigate a significant amount of emissions so that developing regions can mitigate less (as shown in Figure A.9 in Appendix A.8.3), thus improving global social welfare. After the 2030s, however, the US net emissions become zero in the cooperative case [^16] For regional net emissions under cooperation and noncooperation, see Appendix A.8.2. and cannot be further mitigated, so its MAC starts to decline over time. The effect of the ETS under cooperation depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm). Moreover, the ETS has little impact on the net emissions and the MAC of a developed region as shown in Figures A.7 and A.9 (in Appendices A.8.2 and A.8.3) respectively, because a developed region will abate significantly no matter whether the ETS exists or not under cooperation. That is, the ETS plays an important role only in developing regions under cooperation.
@@
The top-right panel of Figure 7 shows that the cooperative SCC of the US is sustained at a high level throughout all periods. This is consistent with the high MAC of the US. When the US regional net emissions are not binding at zero before 2030 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2030. Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions. This also implies that a developed region will contribute more to mitigating emissions under cooperation. This pattern is consistent across regions: Figure A.11 in Appendix A.8.4 shows that the US, EU, Japan and OHI have a high cooperative SCC, while the developing regions like India and Africa have a small cooperative SCC.
Under noncooperation, the MAC of the US is substantially lower in the case with the ETS compared to the case without the ETS (Figure 7, bottom-left panel). Such discrepancy persists until 2058, the period during which the US is a net permit buyer (shown in Figure 3), reflecting the fact that implementation of the ETS offers the US a more cost-effective means of keeping its net emissions below its cap. Consequently, the US conducts less emission abatement. For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2030 (see Figure A.10 in Appendix A.8.3).
@@
Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS. For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2033 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2033 (see the two top panels of Figure 7). For the noncooperative case without the ETS, the US regional net emissions are binding at their caps after 2022 (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC after 2022 (see the two bottom panels of Figure 7). Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Eurasia under cooperation until 2060 in Figures A.7, A.9, A.11; and Eurasia under noncooperation until 2049 in Figures A.8, A.10, and A.12), which is equal to the optimal regional carbon tax. This relationship holds for every region as shown in Figures A.7-A.12.
The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under cooperation or noncooperation. This pattern holds for every region as shown in Figures A.11 and A.12. This happens under noncooperation because the regional noncooperative SCC is the present value of of future climate damages in the region resulting from an additional unit of global emissions released in the current period, and the ETS has little impact on the marginal damages and our endogenous social discount rates. Under cooperation, since the ETS has little impact to developed regions and the temperature path, the wealth redistribution among developing regions via the ETS has little impact on the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$,, so the ETS has little impact on the regional cooperative SCC. This shows that the regional SCC is complementary to the ETS, as the emission caps exist no matter whether the ETS exists or not.
@@ 6.3 Alternative Net Zero Scenarios
Along with the baseline emission cap scenario, we further analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090. The top-left and top-right panels in Figure 8 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,593 per ton of carbon in 2049, and the temperature rise is restricted to 1.61 degrees Celsius by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$445 and \$261 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.8 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
@@
The bottom-left and bottom-right panels in Figure 8 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.8.5, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2085 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$218 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$198 in 2100 in the net zero 2090 scenario. This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.